evaluate ∫ (m/x + x/m + m^x + x^m mx) dx

evaluate
∫ (m/x + x/m + m^x + x^m mx) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int (\frac{m}{x} + \frac{x}{m} + m^{x} + x^{m}.mx)dx
I = \int (\frac{m}{x} + \frac{x}{m} + m^{x} + mx^{m+1})dx
I = \int \frac{m}{x}dx + \int \frac{x}{m}dx + \int m^{x}dx + \int mx^{m+1}dx
I = mln(x) + \frac{x^{2}}{2m} + \frac{m^{x}}{ln(m)} + m.\frac{x^{m+2}}{m+2}+constant